AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 28 of 33  ELOG logo
  ID Date Author Subject
  115   Sat Oct 9 09:47:33 2021 CBTo do list for next week

0. Give split-funnels to OG or Grossmontage to clean. Mount back foot support for gate valves. Critical when we split.

1. Wait for vacuum division to come to install: IE514 cables to current CM52 module, ion pump cables + controls + turn on, RGA head + program + turn on

2. Vacuum division should also bring N2 gas to vent CARME. Ideally vent by Wed to avoid losing time. Remember to disable ion pumps and remove all cables when venting.

3. While waiting to vent, connect pneumatic hoses to motors box in CARME rack. There is one pneumatic in (central feedthrough) and four pneumatic out. See picture attached. Top two go to left pneumatic motor, bottom two go to right pneumatic motor. Of each set of two, one is for extending the motor and one is for retracting but unclear which is which. Try to connect and operate motor manually by pressing on buttons on top of pressure reducer to ensure everything works. Don't forget to remove the safety pin.

4. Some disassembly of the servo motors may be required to get them operational. Use the Windows ClearView program provided by Teknic to control servos and change the function of the limit switches. Download from https://teknic.com/downloads/ -> ClearPath -> Software -> ClearView Install.zip . There is a Linux version but I have never tested it. The code on the raspberry Pi cannot reproduce some of the functions in the official program due to library limitations (but can do other things the official program cannot do).

 

5. After venting CARME, split open at the COF600 flange close to the magnet. Open CARME by 1-2 cm, remove spherical washers from top of rail studs, and lift with crane. Make sure knife edges of the two halves do not touch. CARME currently has feet mounted and cannot easily be placed on a standard pallet. Either dismount the feet (not easy to remount), or find a pallet on which CARME can land on its belly only, or place CARME on pallet with feet and screw feet down? Discuss strategy with DR asap.

6. Install split-funnels on flanges/wire harnesses in use (no point installing on other flanges). This may require dismounting top flanges. Probably switch from 3 hole to 4 hole configuration with funnels.

7. Test detector support going to actuator rod without detector. Make sure you have all the parts, including bolt to secure support to rod. You may want to push the rod in via pneumatic motor (or by hand pushing the plate outside) to gain better access.

8. Take detector support on table, install *two* detectors on same support, and mount on actuator rod in XHV. Connect wires and close via jacking plates. Optionally, consider mouting empty support on other rod? May be useful for diagnostic/scraping?

9. Move XHV thermocouple feedthrough to top flange and connect one thermocouple per detector, one to the support, and leave one floating inside

10. If the motors are fully operational, check how far/close they can be actuated from chamber wall. If not, take some dimensions for future reference.

11. Lift CARME back in place. Close using Pfeiffer COF600 seal and ~72 Nm force per bolt. 80 Nm were required in DL. Probably get Grossmontage to do this. Use our long torque-wrench.

12. Connect to interaction chamber via bellows. Beware gaskets on one side are non standard. One appropriate gasket should be on top of bellows box.

13. Reconnect pumps, rebuild octagon, reconnect thermocouples, rebuild baking tent, bake. Make sure to label the NEG cables this time!

 

14. Activate NEG pumps following this recipe. ALWAYS monitor temperature and be ready to stop if detector temperature reaches 120 C. NEG temperature does not drop immediately, don't wait until it is too late.

If temperature is increasing too quickly (check rates/h), decrease activation power in increments of 25 W. Remember to flash ion pumps.

A. Start from D2000. Activate one at a time with T 400 C, t rise 30 min, t hold 90 min. Activating a second D2000 while the first has reached its hold state seems safe. Leave at 250 C to keep clean.

B. Continue with UVH1400 furthest from the detector (DSSD wall chamber). Activate at P 200 W, t rise 30 min, to hold 90 min. Finish one set of two before moving to the second. Leave at 50 W to keep clean.

C. Continue with UHV1400 in intermediate chamber. Activate at P 150 W, rest as above. Leave at 50 W to keep clean.

D. Continue with 2xUHV1400 closest to detectors. Activate at P 70 W, rest as above. Leave at 50 W to keep clean.

E. Finish with 3xUHV1400 closest to detectors. Activate at P 90 W, rest as above. Turn all NEGs off when activation of 3x is complete.

F. Let chamber cool. Activate ion pumps at room temperature. Section off turbo pump.

  114   Sat Oct 9 09:44:36 2021 CB RSSHeaters removed + motors tests

Removed thermocouples, heaters, RE72 modules and leftover baking tent. Support octagon removed except topmost part which requires disconnecting pneumatic hoses going to top gate valve. To be done on Monday with ML assistance.

Attempted to move servo motors once more - did not work. The motors appear stuck but it is not clear how or where. Removed screws on flanged ball nut and on motor back. Still stuck. Moving by hand is not possible. Code likely complains about excessive force on motor startup. Will discuss with PB. Some dismounting once back in air may be required.

  113   Fri Oct 8 17:56:35 2021 CB RSSMotors test

Attempted first motor test in XHV. Wired and connected limit switches and string potentiometer.

Set up raspberry pi to control motors (.124). Set up pi for remote headless access.

Moved jumper on Teknic hub from 1 to 2 motors configuration.

Linux code only sees one motor and gives issues due to improperly set up limit switches.

Switched to Windows Clear path program. Left servo motor refuses to move and sounds stuck. Unable to move by hand - unclear how braking is still active.

Right servo motor initially moves, but got stuck after one move, on limit switch. Complain about sensorless startup which refers to low bias or excessive torque. The second looks more likely from the display but unclear why. Brake engaged for this motor as well? Unable to move by hand even after disconnecting all cables.

 

Removed safety pin from bellows (compressed air still missing, to be fixed next week). Motors move in. Pressure from 2E-12 to 3.5E-12 mbar and quickly recovers (seconds). Subsequent move cause smaller pressure fluctuation. All good.

 

Will contact PB too figure out issue with motors. Problem unclear. Might require some dismounting.

 

 

 

 

  112   Thu Oct 7 08:40:09 2021 CB RSSCARME in XHV
CARME reached 9E-12 this morning

Ion pumps on
No change in pressure.

Gate valve closed. Turbo sectioned off.

Pressure spikes (normal). Waiting for it 
to recover.

Will try to get rga operational later 
today.
  111   Wed Oct 6 15:36:14 2021 CB RSSNeg activation tests -3
Attempted to activate 2xUHV1400 connected to bottom C4, channel #2 using P 200 W, t rise 30 min, t hold 90 min.
Temperature of Kapton bundle increased drastically to 160+ C. Had to stop. Conclusion is that these are likely
the two UHV1400 on the bottom of CARME, immediately in front of the Kapton bundle.

Left the chamber to cool for a while. Baking temperature set to 85 C.
Repeated attempt using above parameters except P 150 W. Too hot. Had to decrease P to 75 W to remain below 150 C
on the Kapton bundle. Temperature on internals increases slowly, even with P 75 W.
Lesson learned: activate the UHV1400 next to the detectors *last*. Once on, temperature increase on the
detectors can be reduced but not stopped.

Attempted to activate 2xUHV1400 on bottom C4, channel #3 using P 200 W, t rise 30 min, t hold 90 min. These
cannot be close to the detectors.
Possibly due to interference with #2 being on hold, temperature of internals starts increasing. Reduced P to 125
W. OK.
Second lesson learned: temperature on internals is very hard to control once the UHV1400 close to them are
activated. Activate other UHV1400 first.

Set #2 to hold 50 W. Set #3 to hold 50 W.

Attempted to activate *3*xUHV1400 on bottom C4, channel #4 using P 120 W, t rise 30 min, t hold 90 min. These
are definitely close to the detectors.
Temperature increase modest but noticeable. Activation completed before temperature increases
too much.

17:30 German time
Activation complete. All NEGs off simultaneously.
Heating power off.

CARME slowly cools down (<10 C/h)
Opened some parts of the jacket to allow better cooling (still <10 C/h)

Pressure reaches E-11 around 19:00
SIP elements in D2000 still at 120 C. Too hot to turn on.

SIP elements will be turned tomorrow morning. That should conclude this baking cycle.
  110   Wed Oct 6 09:47:30 2021 CB RSSNeg activation tests -2
Attempted activating 2xUHV1400 connected to top C4, channel #3 with P 300 W (450 C?), t rise 30 min, t hold 90
min as yesterday

Power reach 300 W, hold. Temperature increased by 5C. OK.

Activated top C4, channel #4 using same set of parameters. Internal temperature increases significantly, up to
140C and still increasing.

Decreased power on channel #2 to 100 W (down from 200 W yesterday). Should correspond to 350 C or so.
Temperature still increasing.
Decreased power on channels #3 and #4 to 250 W (425 C?). Temperature stops increasing. Around 155C on kapton
bundle. OK for kapton, detectors would have broken.
Baking code detected temperature on internals >150 C and stopped baking for safety. Temperatures do not change -
heating provided by NEGs more than enough to maintain current status.

Decreased power on channel #2 to 50W (250 C?). Temperature on internals decreases slightly. No major effect. 
This is likely because the UHV1400 corresponding to #2 are far from the detectors, while #3 and #4 are closer.

Decided to complete activation of #3 and #4 using 250 W. Set both to 50C after completion. Temperature of
internals decreases quickly.

Re-activated baking code when temperature on all internals <150 C. Setpoint 85C on RE72-1 and 95 on -2.

Now waiting for internals to cool down. Will activate remaining 3 sets (2+2+3 UHV1400) later today with P 200 W
(400 C?) t rise 30 min, t hold 90 min.
We know the set of 3 UHV1400 is definitely close to the kapton bundle. P 300 W (100x3) for that set.
  109   Wed Oct 6 09:39:59 2021 PB CBOriginal dirty assembly of CARME internals
PB found pictures of the original dirty assembly, before screws and components were shipped to DL.
See attached.
  108   Tue Oct 5 17:30:01 2021 CB RSSNEG activation tests
Found D2000 #1 activated and at 250C (25W) 
to prevent getting dirty when other NEGs 
are activated. See previous entries.

Attempted activation of D2000 #2 using T 
390C , rise time 30 m, hold time 90 m

After 30 m temperature inside chamber 
increased by 5C. 390 C corresponds to 75 
W.

Attempted activation of D2000 #3 while #2 
is holding temperature. Same parameters as 
above except T 400.

After 30 min, temperature in chamber up by 
10 more C to 110.

Attempted activation of #4 as above.

Set #2 then #3 then #4 to 250 C after hold 
time is over to protect them. All D2000 
activated. Max temperature on Kapton 
bundle 110 C.



Attempted activation of 2xUHV1400 linked 
to top C4, channel 2.

No thermocouple available. Used power 200 
W (2x100) rise time 30 m and hold time 90 
min. Upon reaching 200 W no effect on 
chamber interior temperature. Increased to 
300 W. Temperatures increased to 120C max 
roughly. Acceptable.

Had to decrease hold time to 60 m to avoid 
remaining in cave alone for too long. 
Likely okay with 300 W (450 C?). Left on 
150W (400C?) to protect from subsequent 
activations tomorrow.

Because cables are unlabeled it's not 
clear how far the activated UHV1400 are 
from the detectors. If close, this recipe 
works. If far, we may have to use lower 
power for other modules.
  107   Tue Oct 5 11:44:11 2021 CB, TD SAES manuals & cable drawings
NEG power
SIP power
UHV1400 operating instructions
  106   Mon Oct 4 20:30:46 2021 RSS JM CB (remotely)CARME baking status -21
Current temperature ~95 C (on manual mode)
See plot and screen attached.

Current pressure 2.86E-10
See
https://uoe-my.sharepoint.com/:x:/g/personal/cbruno2_ed_ac_uk/EVbsVOubNHRLj0PKiFIvd00BClMJ6WSJ1UDAyLA2F3R4jA?rtime=OK4kVZKB2Ug
http://web-docs.gsi.de/~lestinsk/vacplot.php
  105   Mon Oct 4 09:19:11 2021 JM RSSActivation of NEG elements

9:30

Ion pumps de-gassed by switching on and off multiple times. Large pressure increase the first time pumps were flashed, pressure recovers quickly once pumps off. Subsequent flashing of ion pumps showed a significantly smaller pressure increase, unitl little increase was seen. 

see http://web-docs.gsi.de/~lestinsk/vacplot.php

Pressure ~3E-8 mbar

10:00

Top NEG controller activated. This controller controls 6 NEG wafers connected in groups of two. NEG wafers connected in two's have the following settings which were utilised in Daresbury

Ramp Time = 45 minutes

Hold Time  = 60 minutes

Vmax =  100 V

Imax =  5A

Pmax = 400W

(See attachment 4)

Settings for the NEG wafers connected in a group of three (white tape on cable) in the moving detector chamber are the same as above but with Pmax = 600W. (See attachment 3)

Settings for the D2000's (kapton tape on cables) are the same used in Daresbury and are;

Ramp Time = 45 minutes

Hold Time  = 60 minutes

Vmax =  65 V

Imax =  3A

Pmax = 254W

Tmax = 500 C

(See attachment 1 and 2)

11:44 

Activation of first set of NEG's completed. The temperature of internal thermocouples went above 150C during activation so power to NEG elements was reduced during activation to 125W per module. NEG's fully OFF to allow internal temperatures to cool. Modules will be activated alone to try and find way of activating NEG's without high temperatures. 

Current bakeout temperature is 95C

Pressure ~ 8E-9 mbar

13:10

All NEG's off. Bakeout Temperature is 95C, internal temperature are ~102C.

Pressure  = 5.8E-10 mbar

Activating a single D2000 NEG pump as they contain in built thermocouples to gauge the effect on internal temperatures. 

14:53

Activation complete for single D2000 up to 500C. Most internal thermocouples remained <110C during activation with one going up to 140C. Temperature of internal thermocouples largely unaffected by the NEG activation untill above 400C NEG temperature. D2000 element will be maintained at 250C for activation of other elements. 

Internal temperatures fallen back to ~102C

Pressure <4E-10 mbar

Would recommend activating elements one at a time with a 20-30 minute ramp and limiting power for NEG modules to ~100W per module. 

  104   Mon Oct 4 06:34:58 2021 RSS JM CB (remotely)CARME baking status -20
Current temperature ~110 C (on manual mode)
See plot and screen attached.

Current pressure 3.20E-8
See
https://uoe-my.sharepoint.com/:x:/g/personal/cbruno2_ed_ac_uk/EVbsVOubNHRLj0PKiFIvd00BClMJ6WSJ1UDAyLA2F3R4jA?rtime=OK4kVZKB2Ug
http://web-docs.gsi.de/~lestinsk/vacplot.php
  103   Sun Oct 3 08:36:01 2021 RSS JM CB (remotely)CARME baking status -19
08:36
Current temperature ~132 C (on manual mode RE72-1: 132 C, RE72-2: 130 C)
See plot and screen attached.

Current pressure 1.41E-7

19:42
Temperature is being slowly reduced (<10C/h) to ~100C in order to activate NEG pumps Monday morning. 
Current pressure ~1.2E-7 and will begin to drop. 

23:20
Temperature set point 110C. 
Current pressure ~5E-8

See
https://uoe-my.sharepoint.com/:x:/g/personal/cbruno2_ed_ac_uk/EVbsVOubNHRLj0PKiFIvd00BClMJ6WSJ1UDAyLA2F3R4jA?rtime=OK4kVZKB2Ug
http://web-docs.gsi.de/~lestinsk/vacplot.php
  102   Sat Oct 2 22:14:14 2021 RSS JM CB (remotely)CARME baking status -18
Current temperature ~132 C (on manual mode RE72-1: 132 C, RE72-2: 130 C)
See plot and screen attached.

Current pressure 1.60E-7
See
https://uoe-my.sharepoint.com/:x:/g/personal/cbruno2_ed_ac_uk/EVbsVOubNHRLj0PKiFIvd00BClMJ6WSJ1UDAyLA2F3R4jA?rtime=OK4kVZKB2Ug
http://web-docs.gsi.de/~lestinsk/vacplot.php
  101   Sat Oct 2 08:26:49 2021 RSS JM CB (remotely)CARME baking status -17
Current temperature ~132 C (on manual mode RE72-1: 132 C, RE72-2: 130 C)
See plot and screen attached.

Current pressure 1.73E-7
See
https://uoe-my.sharepoint.com/:x:/g/personal/cbruno2_ed_ac_uk/EVbsVOubNHRLj0PKiFIvd00BClMJ6WSJ1UDAyLA2F3R4jA?rtime=OK4kVZKB2Ug
http://web-docs.gsi.de/~lestinsk/vacplot.php
  100   Fri Oct 1 21:47:51 2021 CB (remotely)CARME baking status -16
Current temperature ~132 C (on manual mode RE72-1: 132 C, RE72-2: 130 C)
See plot and screen attached.

Current pressure 2.03E-7
See
https://uoe-my.sharepoint.com/:x:/g/personal/cbruno2_ed_ac_uk/EVbsVOubNHRLj0PKiFIvd00BClMJ6WSJ1UDAyLA2F3R4jA?rtime=OK4kVZKB2Ug
http://web-docs.gsi.de/~lestinsk/vacplot.php
  99   Fri Oct 1 07:17:46 2021 RSS JMCARME baking status -15
Current temperature ~132 C (on manual mode RE72-1: 132 C, RE72-2: 130 C)
See plot and screen attached.

Current pressure ~2.30E-7 mbar
See
https://uoe-my.sharepoint.com/:x:/g/personal/cbruno2_ed_ac_uk/EVbsVOubNHRLj0PKiFIvd00BClMJ6WSJ1UDAyLA2F3R4jA?rtime=OK4kVZKB2Ug
http://web-docs.gsi.de/~lestinsk/vacplot.php
  98   Thu Sep 30 18:16:52 2021 RSS JMCARME baking status -14
Current temperature ~132 C (on manual mode RE72-1: 132 C, RE72-2: 130 C)
See plot and screen attached.

Current pressure ~3.30E-7 mbar
See
https://uoe-my.sharepoint.com/:x:/g/personal/cbruno2_ed_ac_uk/EVbsVOubNHRLj0PKiFIvd00BClMJ6WSJ1UDAyLA2F3R4jA?rtime=OK4kVZKB2Ug
http://web-docs.gsi.de/~lestinsk/vacplot.php
  97   Thu Sep 30 06:29:39 2021 RSS JMCARME baking status -13
Current temperature ~130 C (on manual mode)
See plot and screen attached.

Current pressure ~4.07E-7 mbar
See
https://uoe-my.sharepoint.com/:x:/g/personal/cbruno2_ed_ac_uk/EVbsVOubNHRLj0PKiFIvd00BClMJ6WSJ1UDAyLA2F3R4jA?rtime=OK4kVZKB2Ug
http://web-docs.gsi.de/~lestinsk/vacplot.php
  96   Wed Sep 29 15:34:07 2021 RSS JMCARME baking status -12
Current temperature ~130 C (on manual mode)
See plot and screen attached.

Current pressure ~4.51E-7 mbar
See
https://uoe-my.sharepoint.com/:x:/g/personal/cbruno2_ed_ac_uk/EVbsVOubNHRLj0PKiFIvd00BClMJ6WSJ1UDAyLA2F3R4jA?rtime=OK4kVZKB2Ug
http://web-docs.gsi.de/~lestinsk/vacplot.php
ELOG V3.1.4-unknown